# Learning large softmax mixtures with warm start EM

### v1

@@ 2.1 Preliminaries
In this section we collect background results on exact parameter recovery in finite mixture models. We begin by recalling a fundamental result in Lemma 1. It shows that the mixture model parameters can be uniquely determined from the moments and mixed moments of $\rho$ defined below. The result is constructive, in that it provides explicit parameter expressions as functions of these moments. In the next section we will make use of these expressions for parameter estimation. Results (11) and (13) below can be found in Lindsay (1989), whereas (12) is implicit in Lindsay and Basak (1993), and we derive its explicit form here.
By the modeling assumption, the true parameters ${\bm{{\bm{\theta}}}}_{1}^{*},\dots,{\bm{{\bm{\theta}}}}_{K}^{*}$ are distinct. The arguments presented below rely on the existence of a unit vector $v\in\mathbb{S}^{L-1}$ such that the inner products ${v}^{\top}{\bm{{\bm{\theta}}}}_{1}^{*}$, $\dots$, ${v}^{\top}{\bm{{\bm{\theta}}}}_{K}^{*}$ are all different from each other; this vector is called the primary axis in Lindsay and Basak (1993). The existence of such a vector is guaranteed, since any ${v}$ that is chosen uniformly at random from $\mathbb{S}^{L-1}$, satisfies
${v}^{\top}{\bm{{\bm{\theta}}}}_{i}^{*}\neq{v}^{\top}{\bm{{\bm{\theta}}}}_{j}^{*}\quad\quad\forall i,j\in[K],\,i\neq j$
with probability one.
Fixing a unit vector $v\in\mathbb{R}^{L}$, and fixing an $L\times L$ matrix $R:=(v,w_{2},\ldots,w_{L})$ with orthonormal columns, we let
${\bm{\theta}}_{k}:=R^{\top}{\bm{\theta}}_{k}^{*}.$
In the sequel, we refer to
$(\theta_{11},\theta_{12},\ldots,\theta_{1K})~{}:=~{}(v^{\top}{\bm{\theta}}_{1}^{*},v^{\top}{\bm{\theta}}_{2}^{*},\ldots,v^{\top}{\bm{\theta}}_{K}^{*})$
as first coordinates. Likewise, the $i$-th coordinate of ${\bm{\theta}}_{k}$ will be $\theta_{ik}:=w_{i}^{\top}{\bm{{\bm{\theta}}}}_{k}^{*}$, for $i\in\{2,\ldots,L\}$, and each $k\in[K]$. In this, and the following three subsections, we assume that $R$ has been chosen, and we show how ${\bm{\theta}}_{k}$, $k\in[K]$, can be recovered at the population and sample level. We devote Section 2.5 to the practical construction of $R$. With this notation, we make the following assumption.
Assumption 1.
$\Delta(\theta_{11},\ldots,\theta_{1K}):=\min_{k\neq k^{\prime}}|\theta_{1k}-\theta_{1k^{\prime}}|>0.$ (7)
To state the results of this section, we let $Z\sim\rho$ be a discrete random vector taking values from $\{{\bm{\theta}}_{1},\ldots,{\bm{\theta}}_{K}\}$ with corresponding probabilities in ${\bm{\alpha}}^{*}=(\alpha_{1}^{*},\ldots,\alpha_{K}^{*})^{\top}$. Its first coordinate $Z_{1}$ has the first $12K-$ 1 moments given by
$m_{r}:=\mathbb{E}_{\rho}(Z_{1})^{r}=\sum_{k=1}^{K}\alpha_{k}^{*}\theta_{1k}^{r},\,\quad 0\leq r\leq 2K-1.$ (8)
Similarly, consider the following mixed-moments:
$m_{r1;i}:=\mathbb{E}_{\rho}\left[Z_{1}^{r}Z_{i}\right]=\sum_{k=1}^{K}\alpha_{k}^{*}\theta_{1k}^{r}\theta_{ik},\quad 0\leq r\leq K-1,~{}2\leq i\leq L.$ (9)
@@
If Assumption 1 holds, then $\bm{m}$ and $\bm{m}_{1;i}$ for $2\leq i\leq L$ uniquely determine ${\bm{\alpha}}^{*}\in\Delta_{K}$ and ${{\bm{\theta}}}_{1},\ldots,{{\bm{\theta}}}_{K}\in\mathbb{R}^{L}$, up to some permutation, as shown below.
@@
    The first coordinates $\theta_{11},\theta_{12},\ldots,\theta_{1K}$ are the unique $K$ roots of the degree $K$ polynomial $P(x)$, in one variable, given by
@@
    For each $k\in[K]$, the remaining $L-1$ coordinates $\theta_{ik}$, $i\in\{2,\ldots,L\}$ are uniquely given by
    $\theta_{ik}=\begin{pmatrix}m_{01;i}\\ \vdots\\ m_{(K-1)1;i}\end{pmatrix}^{\top}\begin{pmatrix}1&m_{1}&\dots&m_{K-1}\\ m_{1}&m_{2}&\dots&m_{K}\\ \vdots&\vdots&&\vdots\\ m_{K-1}&m_{K}&\dots&m_{2K-2}\end{pmatrix}^{-1}\begin{pmatrix}1\\ \theta_{1k}\\ \vdots\\ \theta_{1k}^{K-1}\end{pmatrix}.$ (12)
@@
    ${\bm{\alpha}}^{*}=\begin{pmatrix}1&1&\ldots&1\\ \theta_{11}&\theta_{12}&\ldots&\theta_{1K}\\ \vdots&\vdots&&\vdots\\ \theta_{11}^{K-1}&\theta_{12}^{K-1}&\ldots&\theta_{1K}^{K-1}\end{pmatrix}^{-1}\begin{pmatrix}1\\ m_{1}\\ \vdots\\ m_{K-1}\end{pmatrix}.$ (13)
The proof is given, for completeness, in Section A.1.
@@
Note that it is easy to see that $\bm{\alpha}^{*}$ and ${{\bm{\theta}}_{1}},\dots,{{\bm{\theta}}_{K}}$ are uniquely determined given all moments of the form
$m_{r_{1},\dots,r_{L}}:=\mathbb{E}_{\rho}[(\bm{e}_{1}^{\top}Z)^{r_{1}}\cdots(\bm{e}_{L}^{\top}Z)^{r_{L}}]\quad\quad r_{1},\dots,r_{L}\in\mathbb{N}\,,$
@@
where $\bm{e}_{1},\ldots,\bm{e}_{L}$ denote the canonical basis vectors in $\mathbb{R}^{L}$. Knowledge of these moments is equivalent to knowledge of $\mathbb{E}_{\rho}[(v^{\top}Z)^{r}]$ for all $v\in\mathbb{S}^{L-1}$ and $r\in\mathbb{N}$, which determines the measure $\rho$. The virtue of Lemma 1 is that it identifies a minimal collection of moments of this type. Indeed, one can show that a strict subset of $\bm{m}$ and $\bm{m}_{i}$, for $2\leq i\leq L$, fails to uniquely identify $\rho$.
In the next section we show that the moments and mixed-moments of the mixing measure, $\bm{m}$ and $\bm{m}_{1;i}$ of Lemma 1, can be consistently estimated from samples generated as in (1) and (3). The construction of these estimates, given in Section 2.2 below, is one of our main contributions. We then use Lemma 1, with population-level quantities replaced by appropriate estimators, to obtain estimates of the softmax mixture parameters.
For this purpose, in Section 2.2 it is enough to estimate the latent moments when $X_{1},\ldots,X_{p}$ i.i.d. $\mu$, with $\mu=\mathcal{N}_{L}(0,\mathbf{I}_{L})$ being the standard Gaussian density on $\mathbb{R}^{L}$. In Section 2.4 we show that this procedure can be used for parameter estimation when $\mu=\mathcal{N}_{L}(0,\Sigma)$, with arbitrary covariance matrix $\Sigma$, via a simple re-scaling argument. We remark that the zero mean assumption can be made without loss of generality since the softmax function is invariant under translation by the same value in each coordinate.
2.2 Latent moment estimation in softmax mixtures
We construct estimators of the mixing measure moments in two stages. To illustrate the general idea, it is enough to consider estimation of the general $r$-th moment $m_{r}$ given by (8) above.
Recall that we observe a sample $Y_{1},\ldots,Y_{N}$ from the mixture distribution, which is i.i.d. as  $Y\mid X_{1}=x_{1},\ldots,X_{p}=x_{p}\sim\pi_{{\bm{\omega}}^{*}}(\cdot\mid x_{1},\ldots,x_{p})$. Pick any $r\in\mathbb{N}$. Since information about ${\bm{\omega}}^{*}$ is implicit in this sample, we are naturally lead to searching for a function $h_{r}:\mathbb{R}^{L}\to\mathbb{R}$ such that
$\widehat{m}_{r}:=\frac{1}{N}\sum_{\ell=1}^{N}h_{r}(Y_{\ell})$ (14)
can serve as the estimator of $m_{r}$ defined in (8). Should such a function $h_{r}$ be found, we first note that $\widehat{m}_{r}$ will be close to its conditional mean
$\bar{m}_{r}:=\mathbb{E}(\widehat{m}_{r}\mid X_{1},\ldots,X_{p})=\mathbb{E}(h_{r}(Y)\mid X_{1},\ldots,X_{p}).$ (15)
As we ultimately want to estimate $m_{r}$, we thus must construct a function $h_{r}$ such that $\bar{m}_{r}$ is, in turn, close to $m_{r}$.
We make this rationale formal below, and begin by showing the construction of functions $h_{r}$ that meet our goal. We begin by noting that for any $h_{r}:\mathbb{R}^{L}\rightarrow\mathbb{R}$,
$\bar{m}_{r}$ $=\sum_{k=1}^{K}\alpha_{k}^{*}{{1\over p}\sum_{j=1}^{p}h_{r}(X_{j})\exp(X_{j}^{\top}{\bm{\theta}}_{k}^{*})\over{1\over p}\sum_{i=1}^{p}\exp(X_{i}^{\top}{\bm{\theta}}_{k}^{*})},$ (16)
since the expectation in (15) is under the softmax mixture given by (2) and (3). We could view $\bar{m}_{r}$ as a quasi-estimator, as it involves averages over the first sample. As we seek to approximate $m_{r}$ given by (8) by $\bar{m}_{r}$ above, it is therefore enough to construct a function $h_{r}$ such that, for $X\sim\mathcal{N}_{L}(0,\mathbf{I}_{L})$ and a generic ${\bm{\theta}}\in\mathbb{R}^{L}$, we have
${\mathbb{E}\left[h_{r}(X)\exp(X^{\top}{\bm{\theta}})\right]\over\mathbb{E}\left[\exp(X^{\top}{\bm{\theta}})\right]}=(v^{\top}{\bm{\theta}})^{r},\,$
since with ${\bm{\theta}}={\bm{\theta}}_{k}^{*}$ for each $k\in[K]$, the $r$th moment $m_{r}$ in (8) will be the a.s. $\mu$ limit of the ratio in the right hand side of (16).
Implicitly, the construction of a function $h_{r}$ that satisfies the above equality will depend on the given $v$, as seen below, but we omit the explicit dependency to avoid heavy notation.
We will also need to approximate the mixed moments given by (9). Reasoning as above, we will further construct appropriate functions $h_{r1;i}$ such that
$\bar{m}_{r1;i}:=\mathbb{E}(h_{r1;i}(Y)\mid X_{1},\ldots,X_{p})$ (17)
converges to $m_{r1;i}$, and then construct
$\widehat{m}_{r1;i}:=\frac{1}{N}\sum_{\ell=1}^{N}h_{r1;i}(Y_{\ell})$ (18)
as the final estimators of $m_{r1;i}$.
We let the classical probabilist’s Hermite polynomials $\{H_{r}\}_{r\geq 0}$ be defined by
$H_{r}(x)=r!\sum_{m=0}^{\lfloor r/2\rfloor}\frac{(-1)^{m}}{m!(r-2m)!2^{m}}x^{r-2m}.$
In terms of these polynomials, and for a given orthonormal basis $\{v,w_{2},\ldots,w_{L}\}$ as in Section 2.1, we define $h_{r}$ and $h_{r1;i}$ by:
$h_{r}(X)$ $=H_{r}(X^{\top}v),$ (19)
$h_{r1;i}(X)$ $=H_{r}(X^{\top}v)(X^{\top}w_{i})\,.$ (20)
These functions have the desired property, as the next proposition shows.
@@ 2.2 Latent moment estimation in softmax mixtures
Let $h_{r}$ and $h_{r1,i}$ be defined as above. Let $X\sim\mathcal{N}_{L}(0,\mathbf{I}_{L})$. Then for any ${\bm{\theta}}\in\mathbb{R}^{L}$,
$\mathbb{E}\left[h_{r}(X)\exp(X^{\top}{\bm{\theta}})\right]\over\mathbb{E}\left[\exp(X^{\top}{\bm{\theta}})\right]$ $=(v^{\top}{\bm{\theta}})^{r}\,$ (21)
$\mathbb{E}\left[h_{r1;i}(X)\exp(X^{\top}{\bm{\theta}})\right]\over\mathbb{E}\left[\exp(X^{\top}{\bm{\theta}})\right]$ $=(v^{\top}{\bm{\theta}})^{r}(w_{i}^{\top}{\bm{\theta}})\,,$ (22)
for $i\in\{2,\ldots,L\}$.
Remark 2.
The proof of Proposition 1 is given in the Section A.2, and is valid more generally, when $X$ has a strictly positive ${\mathcal{C}}^{\infty}$ density on $\mathbb{R}^{L}$ whose moment generating function is finite everywhere and whose mixed partial derivatives of all orders decay super-exponentially at infinity.
Next, for the estimators given by (14) and (18), we form the vectors
$\widehat{\bm{m}}=\left(\widehat{m}_{1},\ldots,\widehat{m}_{2K-1}\right)^{\top}\quad\text{and}\quad\widehat{\bm{m}}_{1;i}:=\left(\widehat{m}_{01;i},\ldots,\widehat{m}_{(K-1)1;i}\right)^{\top}$
as well as their population level counterparts. The next proposition gives their rates of convergence to the targets $\bm{m}$ and $\bm{m}_{1;i}$, $i\in\{2,\ldots,L\}$. Its proof is stated in Section A.3.
Proposition 2.
Assume that $X_{1},\ldots,X_{p}$ are i.i.d. $\mu=\mathcal{N}_{L}(0,\mathbf{I}_{L})$. Assume that, conditionally on $X_{1},\ldots,X_{p}$, we have $Y_{1},\ldots,Y_{N}\mid X_{1}=x_{1},\ldots,X_{p}=x_{p}$ that are i.i.d. $\pi_{{\bm{\omega}}}(\cdot\mid x_{1},\ldots,x_{p})$ given by (2) and (3). Then, a.s. $\mu$, and for all $p\geq 1$:
$\lim_{C\rightarrow\infty}\limsup_{N\rightarrow\infty}\mathbb{P}\left(\|\bm{m}-\widehat{\bm{m}}\|_{2}\geq C\left\{\frac{1}{\sqrt{N}}+\sqrt{\frac{\log\log p}{p}}\right\}\mid X_{1},\ldots,X_{p}\right)$ $=0,$
$\lim_{C\rightarrow\infty}\limsup_{N\rightarrow\infty}\mathbb{P}\left(\|\bm{m}_{1;i}-\widehat{\bm{m}}_{1;i}\|_{2}\geq C\left\{\frac{1}{\sqrt{N}}+\sqrt{\frac{\log\log p}{p}}\right\}\mid X_{1},\ldots,X_{p}\right)$ $=0$
for $i\in\{2,\ldots,L\}$.
The moment estimators (14) and (18) may not be bona fide moments of any distribution, a fact that may have repercussions on parameter estimation. This is discussed in detail, for general mixtures, in Lindsay (1989), together with potential corrections that may be difficult to implement. An alternative approach, in the univariate case, was developed by Wu and Yang (2020), who proposed to project the estimated moments onto the set $\mathcal{M}$ of valid moments.
@@
There exists some known absolute constant $B<\infty$ such that
$\max_{k\in[K]}\|{\bm{\theta}}_{k}^{*}\|_{2}\leq B.$
@@
2 in conjunction with $X\sim\mathcal{N}_{L}(0,\mathbf{I}_{L})$ ensures that $\mathbb{E}[X_{j}^{\top}{\bm{\theta}}_{k}^{*}]=\mathcal{O}(1)$ for all $j\in[p]$ and $k\in[K]$ so that the probabilities in $\mathsf{softmax}(X_{1}^{\top}{\bm{\theta}}_{k}^{*},\ldots,X_{p}^{\top}{\bm{\theta}}_{k}^{*})$ are not too spiky. This is crucial in order to have the softmax parametrization be meaningful, as pointed out in Arora et al. (2016).
Given a univariate probability measure $\mu$ supported within $[-B,B]$ for some $B>0$, write $M_{k}(\mu)$ for its $k$th moment. The set $\mathcal{M}$ is defined as
:=$\mathcal{M}:=\left\{(M_{1}(\mu),\dots,M_{2K-1}(\mu)):{\rm supp}(\mu)\in[-B,B]\right\}.$ (23)
The projection of $\widehat{\bm{m}}$ onto this space is obtained by solving
$\widetilde{\bm{m}}=\operatorname*{argmin}_{{\bm{u}}\in\mathcal{M}}\|{\bm{u}}-\widehat{\bm{m}}\|_{2}.$ (24)

### v2

@@ 3.1 Preliminaries
In this section we collect background results on population level parameter recovery from latent moments, in finite mixture models. We begin by recalling a fundamental result in Lemma 1. It shows that the mixture model parameters can be uniquely determined from the moments and mixed moments of the mixing measure defined below. The result is constructive, in that it provides explicit parameter expressions as functions of these moments. In the next section we will make use of these expressions for parameter estimation. Results (32) and (34) below can be found in Lindsay (1989), whereas (33) is implicit in Lindsay and Basak (1993), and we derive its explicit form here.
By the modeling assumption, the true parameters ${\bm{{\bm{\theta}}}}_{1}^{*},\dots,{\bm{{\bm{\theta}}}}_{K}^{*}$ are distinct. The arguments presented below rely on the existence of a unit vector $v\in\mathbb{S}^{L-1}$ such that the inner products ${v}^{\top}{\bm{{\bm{\theta}}}}_{1}^{*}$, $\dots$, ${v}^{\top}{\bm{{\bm{\theta}}}}_{K}^{*}$ are all different from each other; this vector is called the primary axis in Lindsay and Basak (1993). The existence of such a vector is guaranteed, as detailed in Section 3.4.1. For ease of presentation, we assume ${\bm{\theta}}_{1}^{*},\ldots,{\bm{\theta}}_{K}^{*}$ are distinct in their first coordinates:
$\Delta(\theta_{11}^{*},\ldots,\theta_{1K}^{*}):=\min_{k\neq k^{\prime}}|\theta_{1k}^{*}-\theta_{1k^{\prime}}^{*}|>0.$ (28)
Let $\rho^{*}:=\sum_{k=1}^{K}\alpha_{k}^{*}\delta_{{\bm{\theta}}_{k}^{*}}$ be the $K$-atomic measure associated with ${\bm{\omega}}^{*}$. As explained below, one can first recover $\theta_{11}^{*},\ldots,\theta_{1K}^{*}$ and then use them to recover the remaining coefficients $\theta_{ik}^{*}$ for $2\leq i\leq L$ and $k\in[K]$, based on certain moments of $\rho^{*}$.
[…]
$m_{r}:=\mathbb{E}_{\rho^{*}}[Z_{1}^{r}]=\sum_{k=1}^{K}\alpha_{k}^{*}(\theta_{1k}^{*})^{r}.$ (29)
Similarly, consider the following mixed-moments: for $0\leq r\leq K-1$ and $2\leq i\leq L$
$m_{r1;i}:=\mathbb{E}_{\rho^{*}}\left[Z_{1}^{r}Z_{i}\right]=\sum_{k=1}^{K}\alpha_{k}^{*}(\theta_{1k}^{*})^{r}\theta_{ik}^{*}.$ (30)
@@
The following lemma shows that ${\bm{\omega}}^{*}$ can be uniquely recovered from the moments in (31).
@@
For any ${\bm{\omega}}^{*}\in\Omega^{*}$ satisfying (28) and $\min_{k}\alpha_{k}^{*}>0$, the system of equations given by (29) and (30) has a unique solution which equals to ${\bm{\omega}}^{*}$, up to label switching. Moreover, the solution can be found explicitly and is given by the expressions below.
@@
    The first coordinates $\theta_{11}^{*},\theta_{12}^{*},\ldots,\theta_{1K}^{*}$ are the unique $K$ roots of the degree $K$ polynomial $P(x)$, in one variable, given by
@@
    For each $k\in[K]$, the remaining $L-1$ coordinates $\{\theta_{ik}^{*}\}_{2\leq i\leq L}$ are uniquely given by
    $\theta_{ik}^{*}=\begin{pmatrix}m_{01;i}\\ \vdots\\ m_{(K-1)1;i}\end{pmatrix}^{\top}\begin{pmatrix}1&m_{1}&\dots&m_{K-1}\\ m_{1}&m_{2}&\dots&m_{K}\\ \vdots&\vdots&&\vdots\\ m_{K-1}&m_{K}&\dots&m_{2K-2}\end{pmatrix}^{-1}\begin{pmatrix}1\\ \theta_{1k}^{*}\\ \vdots\\ (\theta_{1k}^{*})^{K-1}\end{pmatrix}.$ (33)
@@
    ${\bm{\alpha}}^{*}=\begin{pmatrix}1&1&\ldots&1\\ \theta_{11}^{*}&\theta_{12}^{*}&\ldots&\theta_{1K}^{*}\\ \vdots&\vdots&&\vdots\\ (\theta_{11}^{*})^{K-1}&(\theta_{12}^{*})^{K-1}&\ldots&(\theta_{1K}^{*})^{K-1}\end{pmatrix}^{-1}\begin{pmatrix}1\\ m_{1}\\ \vdots\\ m_{K-1}\end{pmatrix}.$ (34)
@@
Note that $\bm{\alpha}^{*}$ and ${{\bm{\theta}}_{1}^{*}},\dots,{{\bm{\theta}}_{K}^{*}}$ are uniquely determined given all moments of the form
$m_{r_{1},\dots,r_{L}}:=\mathbb{E}_{\rho^{*}}[(\bm{e}_{1}^{\top}Z)^{r_{1}}\cdots(\bm{e}_{L}^{\top}Z)^{r_{L}}]\quad\quad\forall~r_{1},\dots,r_{L}\in\mathbb{N}\,,$
@@
where $\bm{e}_{1},\ldots,\bm{e}_{L}$ denote the canonical basis vectors in $\mathbb{R}^{L}$. Knowledge of these moments is equivalent to knowing $\mathbb{E}_{\rho^{*}}[(v^{\top}Z)^{r}]$ for all $v\in\mathbb{S}^{L-1}$ and $r\in\mathbb{N}$, which uniquely determines the measure $\rho^{*}$. The virtue of Lemma 1 lies in identifying a minimal collection of such moments. Indeed, one can show that any strict subset of $\bm{m}$ and $\{\bm{m}_{i}\}_{2\leq i\leq L}$ fails to uniquely identify $\rho^{*}$.
Lemma 1 in fact gives a constructive procedure of recovering ${\bm{\omega}}^{*}$ from appropriate moments of its $K$-atomic measure $\rho^{*}$. When the atomic measure $\rho^{*}$ is interpreted as the mixing measure inducing $\pi(y;{\bm{\omega}}^{*}|x)$, its moments, now viewed as latent, must be estimated from the observable moments of $\pi(y;{\bm{\omega}}^{*}|x)$. The latter is the main novelty in our MoM procedure, and is explained in the next section.
3.2 Population-level global parameter recovery of softmax mixtures with random features
In the following we show that the latent moments and mixed-moments $\bm{m}$ and $\bm{m}_{1;i}$ used by Lemma 1, can be approximated from the moments of $\pi(\cdot;{\bm{\omega}}^{*}|x)=:\pi^{*}(\cdot|x)$. This construction is one of our main contributions. Moreover, it leads to a MoM algorithm that provably yields a parameter $\bar{\bm{\omega}}$ that is within a small neighborhood of ${\bm{\omega}}^{*}$, thereby enabling the EM algorithm, when initialized with this $\bar{\bm{\omega}}$, to recover ${\bm{\omega}}^{*}$ exactly. To illustrate the general idea of approximating the latent moments, it is enough to consider $m_{r}$, given by (29) above, of the first coordinate of $\rho^{*}$. Pick any $r\in\mathbb{N}$. Since information about ${\bm{\omega}}^{*}$ is in $Y\sim\pi^{*}(\cdot|x)$, we are naturally lead to searching for a function $h_{r}:\mathbb{R}^{L}\to\mathbb{R}$ such that
$\bar{m}_{r}:=$ $~\mathbb{E}_{\pi({\bm{\omega}}^{*}\mid x)}[h_{r}(Y)]=~\sum_{k=1}^{K}\alpha_{k}^{*}{{1\over p}\sum_{j=1}^{p}h_{r}(x_{j})\exp(x_{j}^{\top}{\bm{\theta}}_{k}^{*})\over{1\over p}\sum_{\ell=1}^{p}\exp(x_{\ell}^{\top}{\bm{\theta}}_{k}^{*})}$ (35)
is close to $m_{r}=\sum_{k=1}^{K}\alpha_{k}^{*}(\theta_{1k}^{*})^{r}$. Since $x_{1},\ldots,x_{p}$ are i.i.d. realizations from $\mu$, it is therefore enough to construct a function $h_{r}$ such that, for $X\sim\mu$ and a generic ${\bm{\theta}}\in\mathbb{R}^{L}$, we have
${\mathbb{E}_{\mu}\left[h_{r}(X)\exp(X^{\top}{\bm{\theta}})\right]\over\mathbb{E}_{\mu}\left[\exp(X^{\top}{\bm{\theta}})\right]}=\theta_{1}^{r},$ (36)
implying that the right hand side of (35) will be the $\mu$-a.s. limit of $m_{r}$. Similarly, we also need to construct appropriate functions $h_{r1;i}$ for $2\leq i\leq L$, such that
${\mathbb{E}_{\mu}\left[h_{r1;i}(X)\exp(X^{\top}{\bm{\theta}})\right]\over\mathbb{E}_{\mu}\left[\exp(X^{\top}{\bm{\theta}})\right]}=\theta_{1}^{r}\theta_{i}^{r}.$ (37)
This will ensure $\bar{m}_{r1;i}:=\mathbb{E}_{\pi({\bm{\omega}}^{*}|x)}[h_{r1;i}(Y)]$ is close to the mixed-moments $m_{r1;i}$.
It is not clear whether functions $h_{r}$ and $h_{r1;i}$ satisfying (36) and (37) exist for all $\mu$. However, under the following assumption on $\mu$, Proposition 1 below establishes their existence and provides explicit expressions for these functions. Its proof is given in Section C.2. To the best of our knowledge, this is a novel result.
Assumption 2.
@@ 3.2 Population-level global parameter recovery of softmax mixtures with random features
Let $X\sim\mu$ with $\mu$ satisfying 2. For any $r\in\mathbb{N}$ and $i\in\{2,\ldots,L\}$, define, for all $x\in\mathbb{R}^{L}$,
$h_{r}(x)$ $~:=~(-1)^{r}{1\over\mu(x)}~\frac{\mathrm{d}^{r}}{\mathrm{d}t^{r}}\mu(x+t\mathbf{e}_{1})\big{|}_{t=0},$ (38)
$h_{r1;i}(x)$ $~:=~(-1)^{r+1}{1\over\mu(x)}~\frac{\mathrm{d}^{r+1}}{\mathrm{d}t^{r}\mathrm{d}s}\mu(x+t\mathbf{e}_{1}+s\mathbf{e}_{i})\big{|}_{t,s=0}.$ (39)
Then for any given ${\bm{\theta}}\in\mathbb{R}^{L}$, both (36) and (37) hold.
Example 2 (Explicit choices of $h_{r}$).
When $\mu=\mathcal{N}_{L}(a,\Sigma)$, the functions given by (38) or (39) take more familiar forms, and can be expressed in terms of the classical probabilist’s Hermite polynomials $\{H_{r}\}_{r\geq 0}$, defined by
$H_{r}(x)=r!\sum_{b=0}^{\lfloor r/2\rfloor}\frac{(-1)^{b}}{b!(r-2b)!2^{b}}x^{r-2b},\qquad\forall~x\in\mathbb{R}.$ (40)
Then,
$h_{r}(x):=h_{r}(x;a,\Sigma)=\|\Sigma^{-1/2}\mathbf{e}_{1}\|_{2}^{r}H_{r}\left((x-a)^{\top}\Sigma^{-1}\mathbf{e}_{1}/\|\Sigma^{-1/2}\mathbf{e}_{1}\|_{2}\right).$
When $\mu$ is a finite Gaussian mixture $\sum_{j=1}^{J}\lambda_{j}\mathcal{N}_{L}(a_{j},\Sigma_{j})$, then
$h_{r}(x)=\frac{\sum_{j=1}^{J}\lambda_{j}\mu^{(j)}(x)~h_{r}(x;a_{j},\Sigma_{j})}{\sum_{j=1}^{J}\lambda_{j}\mu^{(j)}(x)}.$
Here $\mu^{(j)}$ is the density of $\mathcal{N}_{L}(a_{j},\Sigma_{j})$.
Proposition 1 readily implies that the observable moments $\bar{m}_{r}$ in (35) are close to the true moments $m_{r}$ (and similarly, $\bar{m}_{r1;i}$ to $m_{r1;i}$) in the following sense:
$\bar{m}_{r}-m_{r}$ $=\sum_{k=1}^{K}\alpha_{k}^{*}\left\{{{1\over p}\sum_{i=1}^{p}[h_{r}(x_{i})\exp(x_{i}^{\top}{\bm{\theta}}_{k}^{*})]\over{1\over p}\sum_{j=1}^{p}[\exp(x_{j}^{\top}{\bm{\theta}}_{k}^{*})]}-{\mathbb{E}_{\mu}[h_{r}(X)\exp(X^{\top}{\bm{\theta}}_{k}^{*})]\over\mathbb{E}_{\mu}[\exp(X^{\top}{\bm{\theta}}_{k}^{*})]}\right\}$ (41)
The population-level MoM algorithm thus recovers ${\bm{\omega}}^{*}$ based on applying a variant version of Lemma 1 to $\bar{\bm{m}}=(\bar{m}_{0},\ldots,\bar{m}_{2K-1})^{\top}$ and $\bar{\bm{m}}_{1;i}=(\bar{m}_{01;i},\ldots,\bar{m}_{(K-1)1;i})^{\top}$, as detailed below.
To recover the first coordinates $\theta_{11}^{*},\ldots,\theta_{1K}^{*}$, Lemma 1 requires solving for the $K$ roots of a polynomial that uses $\bar{m}_{r}$ in place of $m_{r}$. This in turn requires the entries of $\bar{\bm{m}}$ to be bona fide moments of a distribution, a condition that is not guaranteed in general. This is discussed in detail, for general mixtures, in Lindsay (1989), together with potential corrections that may be difficult to implement. An alternative approach, in the univariate case, was developed by Wu and Yang (2020), who proposed to project the moments $\bar{\bm{m}}$ onto the set $\mathcal{M}$ of valid moments.
@@
There exists some known constant $B<\infty$ such that $\max_{k\in[K]}\|{\bm{\theta}}_{k}^{*}\|_{2}\leq B.$
@@
3 in conjunction with 2 ensures that $\mathbb{E}_{\mu}[X_{j}^{\top}{\bm{\theta}}_{k}^{*}]=\mathcal{O}(1)$ for all $j\in[p]$ and $k\in[K]$ so that the probabilities in $A({\bm{\theta}}_{k}^{*})$ are not too spiky. This is crucial in order to have the softmax parametrization be meaningful, as pointed out in Arora et al. (2016).
[…]
:=$\mathcal{M}:=\left\{(M_{1}(\nu),\dots,M_{2K-1}(\nu)):{\rm supp}(\nu)\in[-B,B]\right\}.$ (42)
The projection of $\bar{\bm{m}}$ onto this space is obtained by solving
$\widetilde{\bm{m}}=\operatorname*{argmin}_{{\bm{u}}\in\mathcal{M}}\|{\bm{u}}-\bar{\bm{m}}\|_{2}.$ (43)